# Emergent Matryoshka doll-like point gap in a non-Hermitian quasiperiodic lattice

### v1

@@ Abstract
We propose a geometric series modulated non-Hermitian quasiperiodic lattice model, and explore its localization and topological properties. The results show that with the ever-increasing summation terms of the geometric series, multiple mobility edges and non-Hermitian point gaps with high winding number can be induced in the system. The point gap spectrum of the system has a Matryoshka doll-like structure in the complex plane, resulting in a high winding number. In addition, we analyze the limit case of summation of infinite terms. The results show that the mobility edges merge together as only one mobility edge when summation terms are pushed to the limit. Meanwhile, the corresponding point gaps are merged into a ring with winding number equal to one. Through Avila’s global theory, we give an analytical expression for mobility edges in the limit of infinite summation, reconfirming that mobility edges and point gaps do merge and will result in a winding number that is indeed equal to one.
@@ I Introduction
Different from random disorder, quasiperiodic systems, which are intermediate between disorder and order system, are widely used to study localization phase transitions and MEs in low-dimensional systems. A widely known $1D$ quasiperiodic system is called the Aubry-André (AA) model, whose transition point from extended to localized phases can be exactly determined by self-duality conditions SAubry1980 ; SYJitomirskaya1999 . The self-duality condition also predicts that the extended and localized states cannot coexist and that there are no MEs. Consequently, generalized versions of the AA model have been proposed to study MEs, including modulated lattice structures, modified quasiperiodic potentials APadhan2022 ; SDasSarma1988 ; SDasSarma1990 ; TLiu2022 ; SGaneshan2015 ; XLi2017 ; HYao2019 ; XLi2020 ; YCZhang2022 ; XCZhou2023 ; YWang2020 ; XPLi2016 ; ZLu2022 ; SRoy2021 , and the introduction of long-range hopping JBiddle2011 ; XXia2022 ; XDeng2019 ; MGoncalves2023 ; JBiddle2010 . Beyond these, the AA model has found widespread applications in exploring topological Anderson insulators JLi2009 ; GQZhang2021 ; DWZhang2020b ; DWZhang2020c , transport properties YEKraus2012 ; JCCharlier2007 ; SHuang2024 , entanglement transitions LZhou2024 ; LZhou2024a ; LZhou2023 ; SZLi2024 , and multifractal phases XBeng2019 ; MSarkar2022 ; SZLi2023 ; XLin2023 ; HYao2019 ; MGoncalves2023 ; FLiu2015 ; XCZhou2023 . These remarkable properties can be experimentally realized using cold atomic systems FAAn2018 ; DTanese2014 ; GModugno2010 ; MLohse2016 ; SNakajima2016 ; DWZhang2018 , optical lattices FAAn2018 ; HPLuschen2018 ; SAGredeskul1989 ; DNChristodoulides2003 ; YEKraus2012 ; TPertsch2004 ; HPL2018 ; MSchreiber2015 ; MLohse2016 ; SNakajima2016 , photonic crystals YLahini2009 ; MVerbin2013 ; MVerbi2015 ; PRoushan2017 ; PRoushan2017 , and superconducting qubits HLi2023 .
@@
Recently, there has been a growing interest in localization and topological phase transitions in non-Hermitian quasiperiodic systems PHe2022 ; SZLi2024b ; SZLi2024a ; APadhan2024 ; DWZhang2020b ; HJiang2019 ; SLonghi2019 ; YLiu2020 ; QBZeng2020a ; TLiu2020 ; XCai2021 ; YLiu2021a ; JClaes2021 ; LZTang2021 ; SLonghi2021a ; SLonghi2021b ; LZhou2022 ; XCai2022 ; WHan2022 ; QLin2022 ; YLiu2021 ; LJZhai2020 ; LZTang2022 ; TQian2024 ; YLiu2021c ; QBZeng2020b , where non-Hermiticity can stem from non-reciprocal hopping or complex on-site potential SLonghi2019 ; HJiang2019 ; QBZeng2020b . Essentially, the two types of non-Hermitian modulations are related. One can uncover this relation through the dual transformation, that complex on-site potential is equivalent to non-reciprocal hopping in the dual space, vice versa. Furthermore, when there are MEs in the system, the above two non-Hermitian modulation will exhibit the dual relationship. Specifically, under the condition of complex on-site potential (non-reciprocal hopping) modulation, the extended states (localized states) possess the real spectra, while the localized states (extended states) correspond to the complex spectra. Such complex spectra form closed loops in the complex plane and cause the system to exhibit non-zero topological winding numbers ZGong2018 ; KZhang2020 ; PHe2020 . The Anderson transition point in energy coincides with the PT symmetry breaking LWZhou2022 ; SLonghi2019 ; TLiu2022a ; XXia2022 ; SLonghi2021 ; YLiu2021c ; YLiu2021 ; YLiu2020 ; YLiu2021a .
@@
The main findings of this manuscript are as follows. Multiple MEs and Matryoshka doll-like higher-spectrum point gap can emerge in non-Hermitian quasiperiodic lattice system.
The rest of the paper is structured as follows. In Sec. II, we give a brief introduction to the theoretical model and the main findings. In Sec. III, through three cases, we provide the evidence of the emergent Matryoshka doll-like point gaps. By Avila’s global theory, we analytially discuss the case of infinity summation limit in Sec. IV. We summarize this manuscript in Sec. V.
@@ II Model
Let’s start at a non-Hermitian geometric series modulated (GSM) qusiperiodic potential [see Fig. 1(a)]. The corresponding Hamiltonian reads
@@
and $c_{j}^{\dagger}$ ($c_{j}$) denotes the fermion creation (annihilation) operator on site $j$. $J$ represents the nearest-neighboring hopping strength and $L$ is the system size. $V_{j}$ is the GSM potential with quasiperiodic strength $\lambda$. $N$ is the number of summation terms of the geometric series. The corresponding common ratio of the geometric series is $e^{-q}e^{i(2\pi\alpha j+\theta)}$, where $\alpha$ is an irrational number to garantee that the potential energy has a quasiperiodic structure and $\theta$ is the phase offset.
Figure 1: The scheme diagram of GSM lattice model (a) and Matryoshka doll-like point gaps spectra (b).
@@
where $\xi_{m}=\sum_{j=1}^{L}|\psi_{j,m}|^{4}$ is the corresponding inverse participation ratio. $\psi_{j,m}$ denotes the amplitude for the $m$th eigenstate at the $j$th site. For the localized (extended) state, $\Gamma=0$ ($\Gamma=1$), while the critical state corresponds to $\Gamma\in(0,1)$. Note that, $\Gamma$ tending to $1(0)$ would indicate the extended (localized) state at finite size.
@@
which describes the number of times the spectrum revolving around the base point $E_{b}$ in the complex plane ZGong2018 ; SLonghi2019 . Below, we exhibit that Matryoshka doll-like point gaps [see Fig. 1(b)] emerge in the lattice model (1).
III Matryoshka doll-like point gaps
Figure 2: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.2$ (b), $\lambda=0.7$ (c), $\lambda=1$ (d) and $\lambda=1.7$ (e). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=1,~{}1,~{}1,~{}-0.95$ [marked as rhombus in (b)-(e)]. Throughout, we set $L=610$, $\alpha=377/610$ and $N=2$.
@@ III.1 Traditional point gap with $\omega=1$
First, we consider the case of $N=1$. The quasiperiodic potential contains only the first term of the geometric series (2), i.e.,
$V_{j}=\lambda e^{(i2\pi\alpha j+\theta)}.$ (5)
@@ III.2 The point gap with $\omega=2$
Now, let’s turn to point gaps with hight winding numbers. Under the condition of $N=2$, the corresponding quasiperiodic potential reads
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+\lambda e^{-q}e^{i2(2\pi\alpha+j+\theta)}.$ (6)
@@
In Fig. 2(a) we exhibit the fractal dimensions $\Gamma$ versus $\lambda$ and $E$. By contrast to the case of traditional point gap, the Matryoshka doll-like point gaps emerge here accompanied by MEs. In concrete terms, for the former, there is no ME, and only a clean critical point ($\lambda=J$) exists between the localized and extended states. For the latter, however, the potential of one more summation term cause the ME to be in the interval $1.20.48<\lambda<$ 1.2 [see Fig. 2(a)]. Furthermore, the localized states gradually invade the extended region until all eigenstates become localized with an increasing $\lambda$. Meanwhile, as the eigenspectrum changes from pure real to complex, the system’s PT symmetry is gradually broken [see Fig. 2(b)-(e)]. One can also see that the localized state forms two separated circles at the maximum and minimum eigenvalues, which are connected by the eigenvalues of the extended state. And when all the eigenstates are localized, the two circles fuse to form a Matryoshka doll-like structure [see Fig. 2(e)]. Fig. 2(f) exhibits the argument of the determinants $H-E_{b}$ as a function of the flux $\theta$, where the corresponding base points $E_{b}$ are marked as a rhombus in Fig. 2(b)-(e). One can find that the Arg[det($H-E_{b}$)] is always zero with $\theta$ for the extended phase ($\lambda=0.2$), corresponding to the trivial phase with winding number $\omega=0$. Under the condition of $\lambda=0.7$ and $\lambda=1$, the Arg[det($H-E_{b}$)] forms a cycle with the $\theta$, corresponding to the topological phase with winding numbers $\omega=1$. Under the condition of complete localization ($\lambda=1.7$), a topological phase with higher-order winding number $\omega=2$ emerges with an increasing $\theta$.
@@
In other words, two stepwise phase transitions occur in the system, i.e., localization and topological phase transition. These two phase transitions have exactly identical critical points and corresponding regions of different phases. That is to say, when the system gradually changes from the extended phase, through the intermediate phase, and finally into the localized phase, the topological properties of the system also gradually change from the trival phase, through the topological phase of $\omega=1$, and finally into the topological phase of $\omega=2$. Meanwhile, the symmetry of the system will start with PT symmetry, then enter a partial PT symmetry broken region, and finally become a state where PT symmetry is completely broken.
@@
Figure 3: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.8$ (b), $\lambda=1.2$ (c), $\lambda=1.9$ (d). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=0,~{}0,~{}-1.85$ [marked as rhombus in (b)-(d)]. The other parameters $L=610$, $\alpha=377/610$ and $N=3$.
@@ III.3 The point gap with $\omega=3$
The corresponding fractal dimension $\Gamma$ of all eigenstates as a function of $\lambda$ is plotted in the following Fig. 3(a). One can find that there are more MEs and more circles of point gap spectrum [see Fig. 3(b)-(d)]. Furthermore, the point gap’s circles emerge one by one in the complex plane. To be more specific, two circles will appear first at either end of the real energy, and then the third circle structure appears from the middle of the energy spectrum [see Fig. 3(a)]. These circle structures with complex values correspond to localized eigenstates, which are connected together in the complex plane by the pure real eigenspectrum of extended states. Later, with the increase of disorder strength $\lambda$, the three discrete circles will become increasingly larger. Finally, they merge together to form a three-layer Matryoshka doll-like point gap [see Fig. 3(b)-(d)]. The corresponding winding numbers will gradually complete the transformation of $0\rightarrow 1\rightarrow 2\rightarrow 3$. Again, accompanied by topological and localization phase transitions, the PT symmetry gradually breaks. The corresponding arguments confirms the conclusion as shown in Fig. 3(e).
@@ IV Theoretical analysis in infinite limit ($N={\infty}$)
Note that, the expression Eq. (8) is a non-Hermitian version Ganeshan-Pixley-Das Sarma potential SGaneshan2015 , and the corresponding MEs can likewise be solved exactly by the Avila global theory.
@@
and $\gamma_{B}(E)$ can be calculated by Avila global theroy AAvila2015 . We insert an additional imaginary phase $\theta\rightarrow\theta+i\epsilon$. Although $T_{j}\in SL(2,\mathbb{C})$ is not an even function for $\epsilon$, we can handle it by complexification as well YCZhang2022 ; YWang2021 ; YWang2020a ; YWang2021a ; AAvila2015 . Then, by letting $\epsilon\rightarrow-\infty$, one can get
@@
The ciritical energy of MEs satisfies $\gamma=0$. Then, one can obtain
@@
The numerical fractal dimension $\Gamma$ as a function of $\lambda$ are plotted in Fig. 4(a), where the red dashed line is the exact ME $E_{c}$ from the expression Eq. (16). The numerical results are in good agreement with analytical expression.
@@
Furthermore, we can fix the base points $E_{b}=\pm 2$ and compute the winding number $\omega$ as a function of $\lambda$ to characterize the appearance of MEs and the complete localization transition, respectively [see the insets of Fig. 4(b)]. One can see that when the MEs appear, the $\omega$ for the base $E_{b}=-2$ transforms from 0 to 1 and the system has a topological point gap. When $\omega$ for the base point $E_{b}=\pm 2$ are both non-zero, the system enters the localized phase completely. Specifically, we show in the inset the fractal dimension $\Gamma$ in the complex plane under $\lambda=1$ (coexistence of extended and localized states) and $\lambda=4$ (localized phase), which agree with our previous analysis well. Therefore, for the case of $N=\infty$, the localized region encroaches on the extended region from the maximum real energy, and there is always only one topological gap in the system. The corresponding maximum winding number $\omega=1$ (see Appendix B for details about phase diagrams versus $N$).
@@
Figure 4: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane, red dashed line denotes the analytical ME. (b) The winding number versus $\lambda$ for $E_{b}=-2$ (red line) and $E_{b}=2$ (blue line). The insets exhibit point gaps in the complex plane, where the base point $E_{b}=-2,~{}2$ [the rhombus in the insets]. Parameters $L=610$, $\alpha=377/610$ and $N=\infty$.
V Conclusion

### v2

# Emergent multi-loop nested point gap in a non-Hermitian quasiperiodic lattice

@@ Abstract
We propose a geometric series modulated non-Hermitian quasiperiodic lattice model, and explore its localization and topological properties. The results show that with the ever-increasing summation terms of the geometric series, multiple mobility edges and non-Hermitian point gaps with high winding number can be induced in the system. The point gap spectrum of the system has a multi-loop nested structure in the complex plane, resulting in a high winding number. In addition, we analyze the limit case of summation of infinite terms. The results show that the mobility edges merge together as only one mobility edge when summation terms are pushed to the limit. Meanwhile, the corresponding point gaps are merged into a ring with winding number equal to one. Through Avila’s global theory, we give an analytical expression for mobility edges in the limit of infinite summation, reconfirming that mobility edges and point gaps do merge and will result in a winding number that is indeed equal to one.
@@ I Introduction
Different from random disorder, quasiperiodic systems, which are intermediate between disorder and order system, are widely used to study localization phase transitions and MEs in low-dimensional systems. A widely known 1D quasiperiodic system is called the Aubry-André (AA) model, whose transition point from extended to localized phases can be exactly determined by self-duality conditions SAubry1980 ; SYJitomirskaya1999 . The self-duality condition also predicts that the extended and localized states cannot coexist and that there are no MEs. Consequently, generalized versions of the AA model have been proposed to study MEs, including modulated lattice structures, modified quasiperiodic potentials APadhan2022 ; SDasSarma1988 ; SDasSarma1990 ; TLiu2022 ; SGaneshan2015 ; XLi2017 ; HYao2019 ; XLi2020 ; YCZhang2022 ; XCZhou2023 ; YWang2020 ; XPLi2016 ; ZLu2022 ; SRoy2021 , and the introduction of long-range hopping JBiddle2011 ; XXia2022 ; XDeng2019 ; MGoncalves2023 ; JBiddle2010 . Beyond these, the AA model has found widespread applications in exploring topological Anderson insulators JLi2009 ; GQZhang2021 ; DWZhang2020b ; DWZhang2020c , transport properties YEKraus2012 ; JCCharlier2007 ; SHuang2024 , entanglement transitions LZhou2024 ; LZhou2024a ; LZhou2023 ; SZLi2024 , and multifractal phases XBeng2019 ; MSarkar2022 ; SZLi2023 ; XLin2023 ; HYao2019 ; MGoncalves2023 ; FLiu2015 ; XCZhou2023 . These remarkable properties can be experimentally realized using cold atomic systems FAAn2018 ; DTanese2014 ; GModugno2010 ; MLohse2016 ; SNakajima2016 ; DWZhang2018 , optical lattices FAAn2018 ; HPLuschen2018 ; SAGredeskul1989 ; DNChristodoulides2003 ; YEKraus2012 ; TPertsch2004 ; HPL2018 ; MSchreiber2015 ; MLohse2016 ; SNakajima2016 , photonic crystals YLahini2009 ; MVerbin2013 ; MVerbi2015 , and superconducting qubits PRoushan2017 ; HLi2023 .
@@
Recently, there has been a growing interest in localization and topological phase transitions in non-Hermitian quasiperiodic systems PHe2022 ; DSBorgnia2023 ; DSBorgnia2023a ; SZLi2024b ; SZLi2024a ; APadhan2024 ; DWZhang2020b ; HJiang2019 ; SLonghi2019 ; YLiu2020 ; QBZeng2020a ; TLiu2020 ; XCai2021 ; YLiu2021a ; JClaes2021 ; LZTang2021 ; SLonghi2021a ; SLonghi2021 ; LZhou2022 ; XCai2022 ; WHan2022 ; QLin2022 ; YLiu2021 ; LJZhai2020 ; LZTang2022 ; TQian2024 ; YLiu2021c ; QBZeng2020b , where non-Hermiticity can stem from non-reciprocal hopping or complex on-site potential. Essentially, the two types of non-Hermitian modulations are related. One can uncover this relation through the dual transformation, that complex on-site potential is equivalent to non-reciprocal hopping in the dual space, vice versa. Furthermore, when there are MEs in the system, the above two non-Hermitian modulation will exhibit the dual relationship. Specifically, under the condition of complex on-site potential (non-reciprocal hopping) modulation, the extended states (localized states) possess the real spectra, while the localized states (extended states) correspond to the complex spectra. Such complex spectra form closed loops in the complex plane and cause the system to exhibit non-zero topological winding numbers ZGong2018 ; KZhang2020 ; PHe2020 . The Anderson transition point in energy coincides with the PT symmetry breaking LWZhou2022 ; SLonghi2019 ; TLiu2022a ; XXia2022 ; SLonghi2021 ; YLiu2021c ; YLiu2021 ; YLiu2020 ; YLiu2021a .
@@
The main findings of this manuscript are as follows. Multiple MEs and multi-loop nested higher-spectrum point gap can emerge in non-Hermitian quasiperiodic lattice system.
The rest of the paper is structured as follows. In Sec. II, we give a brief introduction to the theoretical model and the main findings. In Sec. III, through three cases, we provide the evidence of the emergent multi-loop nested point gaps. By Avila’s global theory, we analytically discuss the case of infinity summation limit in Sec. IV. In Sec. V, we discuss the physical meaning of non-Hermitian point gap topologies. We summarize this manuscript in Sec. VI.
@@ II Model
Let’s start at a non-Hermitian geometric series modulated (GSM) quasiperiodic potential [see Fig. 1(a)]. The corresponding Hamiltonian reads
@@
and $c_{j}^{\dagger}$ ($c_{j}$) denotes the fermion creation (annihilation) operator on site $j$. $J$ represents the nearest-neighboring hopping strength and $L$ is the system size. $V_{j}$ is the GSM potential with quasiperiodic strength $\lambda$. $N$ is the number of summation terms of the geometric series. The corresponding common ratio of the geometric series is $e^{-q}e^{i(2\pi\alpha j+\theta)}$, where $\alpha$ is an irrational number to guarantee that the potential energy has a quasiperiodic structure and $\theta$ is the phase offset.
Figure 1: The scheme diagram of GSM lattice model (a) and multi-loop nested point gaps spectra (b).
@@
where $\xi_{m}=\sum_{j=1}^{L}|\psi_{j,m}|^{4}$ is the corresponding inverse participation ratio. $\psi_{j,m}$ denotes the amplitude for the $m$-th eigenstate at the $j$-th site. For the localized (extended) state, $\Gamma=0$ ($\Gamma=1$), while the critical state corresponds to $\Gamma\in(0,1)$. Note that, $\Gamma$ tending to $1(0)$ would indicate the extended (localized) state at finite size.
@@
which describes the number of times the spectrum revolving around the base point $E_{b}$ in the complex plane ZGong2018 ; SLonghi2019 . Below, we exhibit that multi-loop nested point gaps [see Fig. 1(b)] emerge in the lattice model (1).
III Multi-loop nested point gaps
Figure 2: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.2$ (b), $\lambda=0.7$ (c), $\lambda=1$ (d) and $\lambda=1.7$ (e). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=1,~{}1,~{}1,~{}-0.95$ [marked as rhombus in (b)-(e)]. Throughout, we set $L=610$, $q=0.2$$\alpha=377/610$ and $N=2$.
@@ III.1 Traditional point gap with $\omega=1$
First, we consider the case of $N=1$. The quasiperiodic potential contains only the first term of the geometric series (2), that is,
$V_{j}=\lambda e^{i(2\pi\alpha j+\theta)}.$ (5)
@@ III.2 The point gap with $\omega=2$
Now, let’s turn to point gaps with high winding numbers. Under the condition of $N=2$, the corresponding quasiperiodic potential reads
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+\lambda e^{-q}e^{2i(2{\pi}{\alpha}{j}+\theta)}.$ (6)
@@
In Fig. 2(a) we exhibit the fractal dimensions $\Gamma$ versus $\lambda$ and $E$. By contrast to the case of traditional point gap, the multi-loop nested point gaps emerge here accompanied by MEs. In concrete terms, for the former, there is no ME, and only a clean critical point ($\lambda=J$) exists between the localized and extended states. For the latter, however, the potential of one more summation term cause the ME to be in the interval $1.20.48<\lambda<$ 1.2 [see Fig. 2(a)]. Furthermore, the localized states gradually invade the extended region until all eigenstates become localized with an increasing $\lambda$. Meanwhile, as the eigenspectrum changes from pure real to complex, the system’s PT symmetry is gradually broken [see Fig. 2(b)-(e)]. One can also see that the localized state forms two separated circles at the maximum and minimum eigenvalues, which are connected by the eigenvalues of the extended state. And when all the eigenstates are localized, the two circles fuse to form a multi-loop nested structure [see Fig. 2(e)]. Fig. 2(f) exhibits the argument of the determinants $H-E_{b}$ as a function of the flux $\theta$, where the corresponding base points $E_{b}$ are marked as a rhombus in Fig. 2(b)-(e). One can find that the Arg[det($H-E_{b}$)] is always zero with $\theta$ for the extended phase ($\lambda=0.2$), corresponding to the trivial phase with winding number $\omega=0$. Under the condition of $\lambda=0.7$ and $\lambda=1$, the Arg[det($H-E_{b}$)] forms a cycle with the $\theta$, corresponding to the topological phase with winding numbers $\omega=1$. Under the condition of complete localization ($\lambda=1.7$), a topological phase with high winding number $\omega=2$ emerges with an increasing $\theta$.
@@
In other words, two stepwise phase transitions occur in the system, i.e., localization and topological phase transition. These two phase transitions have exactly identical critical points and corresponding regions of different phases. That is to say, when the system gradually changes from the extended phase, through the intermediate phase, and finally into the localized phase, the topological properties of the system also gradually change from the trivial phase, through the topological phase of $\omega=1$, and finally into the topological phase of $\omega=2$. Meanwhile, the symmetry of the system will start with PT symmetry, then enter a partial PT symmetry broken region, and finally become a state where PT symmetry is completely broken.
@@
Figure 3: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.8$ (b), $\lambda=1.2$ (c), $\lambda=1.9$ (d). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=0,~{}0,~{}-1.85$ [marked as rhombus in (b)-(d)]. The other parameters $L=610$, $\alpha=377/610$, $q=0.2$ and $N=3$.
@@ III.3 The point gap with $\omega=3$
The corresponding fractal dimension $\Gamma$ of all eigenstates as a function of $\lambda$ is plotted in the following Fig. 3(a). One can find that there are more MEs and more circles of point gap spectrum [see Fig. 3(b)-(d)]. Furthermore, the point gap’s circles emerge one by one in the complex plane. To be more specific, two circles will appear first at either end of the real energy, and then the third circle structure appears from the middle of the energy spectrum [see Fig. 3(a)]. These circle structures with complex values correspond to localized eigenstates, which are connected together in the complex plane by the pure real eigenspectrum of extended states. Later, with the increase of disorder strength $\lambda$, the three discrete circles will become increasingly larger. Finally, they merge together to form a three-layer multi-loop nested point gap [see Fig. 3(b)-(d)]. The corresponding winding numbers will gradually complete the transformation of $0\rightarrow 1\rightarrow 2\rightarrow 3$. Again, accompanied by topological and localization phase transitions, the PT symmetry gradually breaks. The corresponding arguments confirms the conclusion as shown in Fig. 3(e).
@@ IV Theoretical analysis in infinite limit ($N={\infty}$)
Note that, the expression Eq. (8) is a non-Hermitian version of the Ganeshan-Pixley-Das Sarma potential SGaneshan2015 , and the corresponding MEs can likewise be solved exactly by the Avila global theory.
@@
and $\gamma_{B}(E)$ can be calculated by Avila global theory AAvila2015 . We insert an additional imaginary phase $\theta\rightarrow\theta+i\epsilon$. Although $T_{j}\in SL(2,\mathbb{C})$ is not an even function for $\epsilon$, we can handle it by complexification as well YCZhang2022 ; YWang2021 ; YWang2020a ; YWang2021a ; AAvila2015 . Then, by letting $\epsilon\rightarrow-\infty$, one can get
@@
The critical energy of MEs satisfies $\gamma=0$. Then, one can obtain
@@
The numerical fractal dimension $\Gamma$ as a function of $\lambda$ is plotted in Fig. 4(a), where the red dashed line is the exact ME $E_{c}$ from the expression Eq. (16). The numerical results are in good agreement with analytical expression.
@@
Furthermore, we can fix the base points $E_{b}=\pm 2$ and compute the winding number $\omega$ as a function of $\lambda$ to characterize the appearance of MEs and the complete localization transition, respectively [see the insets of Fig. 4(b)]. One can see that when the MEs appear, the $\omega$ for the base $E_{b}=-2$ transforms from 0 to 1 and the system has a topological point gap. When $\omega$ for the base point $E_{b}=\pm 2$ are both non-zero, the system enters the localized phase completely. Specifically, we show in the inset the fractal dimension $\Gamma$ in the complex plane under $\lambda=1$ (coexistence of extended and localized states) and $\lambda=4$ (localized phase), which agree with our previous analysis well. Therefore, for the case of $N=\infty$, the localized region encroaches on the extended region from the maximum real energy, and there is always only one topological gap in the system. The corresponding maximum winding number $\omega=1$ (see Appendix  C for details about phase diagrams versus $N$).
@@
Figure 4: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane, red dashed line denotes the analytical ME. (b) The winding number versus $\lambda$ for $E_{b}=-2$ (red line) and $E_{b}=2$ (blue line). The insets exhibit point gaps in the complex plane, where the base point $E_{b}=-2,~{}2$ [the rhombus in the insets]. Parameters $L=610$, $\alpha=377/610$, $q=0.2$ and $N=\infty$.
V The point gap in dual space
From the previous section, one can find that the emergence of winding numbers and multi-loop nested point gaps is due to the superposition of different terms of the GSM potential. The analysis in dual space, one can further understand the physical meaning of point gaps.
By applying the dual transformation, one can get the dual space Hamiltonian, i.e.,
$H_{k}=\sum_{n=1}^{N}\sum_{k=1}^{L-n}{\lambda}e^{-(n-1){q}}c_{k}^{\dagger}c_{k+n}+\sum_{k=1}^{L}2J\cos(2{\pi}{\alpha}k)c_{k}^{\dagger}c_{k},$ (17)
where a set of unidirectional long-range hopping terms emerge in the expression (17). The parameter $q$ represents the decay strength of hopping terms, and the parameter $N$ determines both the number of hopping terms and the limit distance of the long-range hopping. We plot the dual space phase diagram in the $\lambda-E$ plane based on the fractal dimension property [see Fig. 5(a)]. The blue (yellow) region corresponds exactly to the localized (extended) phase, which means that the localization property in the dual space is opposite to that in lattice space. Furthermore, we exhibit the eigenenergy with $\lambda=1.7$ under PBCs [see Fig. 5(b)]. Consistent with the above conclusions, the eigenstates in the dual space in this case behave as extended states, which are exactly the opposite of the localized states in the lattice space.
On the other hand, directly from the mathematical structure of the dual space Hamiltonian (17), one can find that there exist nonreciprocal hopping terms, which will lead to the emergence of skin effects in the dual space. This is another visual angle for topological point gaps.
Figure 5: (a) The dual space phase diagram. Eigenenergies in the complex plane with $\lambda=1.7$ under PBCs (b) and OBCs (c). (d) The dual space wave function of the eigenstate corresponding to an eigenenergy of $-0.9217$. The other parameters $L=610$, $\alpha=377/610$, $q=0.2$ and $N=2$.
Furthermore, we calculate the eigenspectrum under OBCs [see Fig. 5(c)]. The eigenspectrum completely collapses into the real axis, i.e., all the eigenvalues in the spectrum are pure real. The corresponding fractal dimensions reveal that all the states of the system are localized. As the skin mode, the corresponding wave function exhibits the characteristic of a left skin mode [see Fig. 5(d)]. In other words, the spectral winding number of the model can be used to describe the skin effect in the dual space.
VI Conclusion